OpenStudy (secret-ninja):

Solve for x: -3|x - 3| = -6 x = 1, x = -1 x = 1, x = 5 x = 0, x = 5 No solutions

OpenStudy (secret-ninja):

I am stuck between B and C.

OpenStudy (anonymous):

One of them are right what would you do first to isolate the absolute value

OpenStudy (secret-ninja):

I made the two equations: -3|x-3| = -6 and -3|x+3| = -6 For the first one I solved and got x=5 so I figured it would be B or C but for the second equation I got x=2. Where did I go wrong?

OpenStudy (anonymous):

divide by -3 first x-3=2 x-3=-2

OpenStudy (anonymous):

There is a minus on the whole brackets, not just 3

OpenStudy (anonymous):

-3(-(x-3))=-6

OpenStudy (anonymous):

now isolate x

OpenStudy (anonymous):

-3(-x+3)=-6

OpenStudy (kaederfds):

x can be 1 or 5

OpenStudy (secret-ninja):

-3(x+3) = -6 -3x + 0 = -6 -3x = -6 x=2 That is what I did...

OpenStudy (anonymous):

-3 * 3 = -9

OpenStudy (anonymous):

you need to isolate the absolute value before you do anything -3|x-3|=-6 x-3=2 first equation x-3=-2 second equation

OpenStudy (secret-ninja):

Oh I see what I did... I didn't distribute properly and got zero where the -9 should go. I just resolved and got x=-1.

OpenStudy (secret-ninja):

oh okay

OpenStudy (anonymous):

you shouldnt be getting -9

OpenStudy (secret-ninja):

Okay wait I did it again I got x=1

OpenStudy (secret-ninja):

Is that right? I think its B

OpenStudy (anonymous):

you dont distribute the -3 you have to isolate the absolute value to do this divide itself and -6 by -3

OpenStudy (anonymous):

yes

OpenStudy (secret-ninja):

Oh... I didn't know what to do so I changed the lines to parentheses... I'm not a very smart person XD

OpenStudy (secret-ninja):

Well thanks you all for your help. I want to give you all medals lol.

OpenStudy (anonymous):

no lol you will get 2 equations with the lines isolate the absolute value one equation the solution will be positive the other will be negative